INTERACTION OF VIBRATIONAL MODES IN 1,4-FLUOROBENZYL ALCOHOL AS REVEALED BY RO-TATIONALLY RESOLVED ELECTRONIC SPECTROSCOPY

\underline{ALEXEI E. NIKOLAEV} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260^{a}.

1,4-Fluorobenzyl alcohol (FBA) exhibits a number of closely spaced (∼ 50 cm^{−1}) bands in its vibrationally resolved S_{1} ←S_{0} electronic spectrum at ∼ 269 nm, previously attributed to a -CH_{2}OH torsional mode. We report here a rotationally resolved study of the four members of this progression. All bands can be fit with rigid rotor Hamiltonians; a comparison of the parameters obtained from these fits shows that all bands belong to the same conformer. All bands exhibit a ∼ 300 MHz splitting due to the torsion-rotation interaction. Band 3 and 4 also exhibit splitting (of ∼100 and ∼800 Mhz, respectively) due to interactions with other low energy vibrational modes. The nature of these modes and their possible participation in the energy relaxation pathways will be discussed.

aWork supported by NSF.